ELECTRONIC TRANSITIONS OF YTTRIUM MONOXIDE

Y.W. NG, NA WANG, ANDREW B. CLARK and \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong..

The electronic transition spectrum of yttrium monoxide (YO) in the spectral region between 284nm and 307nm has been recorded using laser ablation/reaction free-jet expansion and laser induced fluorescence (LIF) spectroscopy. The YO molecule was produced by reacting laser-ablated yttrium atoms with O_{2} seeded in argon. Twenty transition bands were observed in that region and a few bands were selected for further study using optical-optical double resonance (OODR) spectroscopy. The excited C^{2}Π state has been reached via the intermediate B^{2}Σ^{+} state from the ground X^{2}Σ^{+} state. The excited sub-states observed so far have Ω = 0.5 and 1.5. A least squares fit of the measured rotational lines yielded molecular constants for the newly observed excited states.